### v1

# Visual Instruction Bottleneck Tuning

Abstract
Despite widespread adoption, multimodal large language models (MLLMs) suffer performance degradation when encountering unfamiliar queries under distribution shifts. Existing methods to improve MLLM generalization typically require either more instruction data or larger advanced model architectures, both of which incur non-trivial human labor or computational costs. In this work, we take an alternative approach to enhance the robustness of MLLMs under distribution shifts, from a representation learning perspective. Inspired by information bottleneck (IB) principle, we derive a variational lower bound of the IB for MLLMs and devise a practical implementation, Visual Instruction Bottleneck Tuning (Vittle). We then provide a theoretical justification of Vittle by revealing its connection to an information-theoretic robustness metric of MLLM. Empirical validation of three MLLMs on open-ended and closed-form question answering and object hallucination detection tasks over 45 datasets, including 30 shift scenarios, demonstrates that Vittle consistently improves the MLLM’s robustness under shifts by pursuing the learning of a minimal sufficient representation.
1 Introduction
In intensive races on the track of frontier-level AI models, we have observed unprecedented achievements through the form of a general-purpose chat assistant known as multimodal large language models (MLLMs) xai_grok3_2025 ; openai_o3_o4mini_2025 ; googlecloud_gemini2_5_pro_2025 that combine a visual encoder with a large language model. Their universal yet flexible question-answering interface enables MLLMs to easily permeate our lives from general problem-solving liang2024survey ; yang2024large to practical applications AlSaad2024vy ; li2024large ; caffagni2024revolution ; Cui_2024_WACV . While these models may achieve human-like or even surpass human-level performance on certain tasks, a critical gap remains in their robustness—particularly in handling input variations that humans process effortlessly.
Human intelligence thrives on the ability to distill a large amount of sensory and cognitive inputs into concise abstract representations, a process akin to conceptual compression turner2006art ; gray2007abstraction . By prioritizing sparse salient features while discarding redundancy, humans can shape a robust prototypical representation of complex data instances that captures a proper level of invariance to low-level superficial features for generalization, yet maintains sensitivity to high-level abstract features for discrimination miller1956magical ; rosch1975cognitive ; zhaoping2025vision . Unfortunately, there are consistent reports implying that the current MLLMs still lag far behind this desired trade-off between invariance and sensitivity zhang2024out ; han2024well ; ye2025beaf ; oh2025understanding .
Specifically, MLLMs fail to produce relevant responses under query distribution shifts. That is, they are vulnerable to processing subtly perturbed samples and long-tail samples oh2025understanding . This limitation partially stems from the difficulty of acquiring diverse high-quality multimodal instruction data at scale. When trained via standard maximum likelihood estimation on this relatively limited amount of instruction data, MLLM tends to fit to data-specific patterns and result in a brittle solution geirhos2020shortcut ; ye2024mmspubench ; liang2025aligning . To enhance generalization, existing efforts typically fall into two categories (1) data-centric approaches, which collect more instruction data zhao2023svit ; li2024llava ; gu2024infinity and processes input in a finer granularity liu2024llavanext ; shen2025long , and (2) model-centric approaches, which scale up the underlying model using more expressive or specialized backbones chen2024expanding ; tong2024cambrian ; shi2025eagle ; bai2025qwen2 . However, both data scaling and model scaling are resource-intensive—requiring significant annotation or computational cost.
In this work, we propose a new approach from a representation-centric view to improve the robustness of MLLMs under distribution shifts. Rather than scaling data or model, we introduce a lightweight, theoretically grounded module that enhances the internal representations of MLLMs via the information bottleneck (IB) principle. While the IB framework has been explored in small-scale or classification settings alemi2017deep ; 8437679 ; wu2020graph ; mahabadi2021variational ; li2025calibrating , integrating it to autoregressive multimodal instruction tuning poses unique challenges due to the complexity of modeling mutual information across high-dimensional, sequential, and heterogeneous modalities. We overcome these barriers by formulating a novel variational lower bound of the IB objective specifically tailored to the multimodal and sequential nature of MLLMs. We further instantiate this formulation as a modular and scalable implementation—Visual Instruction Bottleneck Tuning (Vittle), which inserts one simple bottleneck layer within the LLM backbone. Vittle pursues minimal sufficient representations cover1999elements that try to preserve only response-relevant information while discarding non-essential residual features. To our knowledge, this is the first work to investigate the IB framework for end-to-end instruction tuning of multimodal LLMs, offering a model-agnostic pathway toward building more robust AI systems.
We conduct an extensive evaluation of Vittle across a wide spectrum of multimodal benchmarks to assess its robustness and generalization under distribution shift. Our experiments span 30 distribution shifts covering diverse forms of perturbation (in both vision and language) and long-tail distributions. Through these evaluations, we demonstrate that Vittle consistently improves robustness over standard instruction tuning baselines, without sacrificing performance on standard benchmarks and canonical tasks. Notably, we find that the bottlenecked representations induced by Vittle lead to enhanced invariance in the latent space, aligning semantically similar inputs more closely—even under input shifts—while reducing overfitting to modality-specific artifacts. We also show that Vittle is compatible with different MLLMs, offering robustness gains while maintaining similar inference-time cost. These results underscore the practical benefit and theoretical promise of information-regularized representation learning for robust multimodal instruction tuning.
Contributions: (1) We propose a new representation-centric framework for improving the robustness of MLLMs under distribution shifts, grounded in the information bottleneck principle. (2) We explore the IB-based end-to-end learning objective of an MLLM for the first time by inducing a new variational lower bound of IB for MLLM and devising a practical instantiation, Vittle, supported by theoretical analysis. (3) Through experiments on 30 diverse types of distribution shifts, we thoroughly validate the robustness of MLLMs on open-ended/closed-form QA and object hallucination detection tasks and show advantages of compressive representation induced by pursuing the IB principle.
2 Background, Related Work, and Motivation
Figure 1: Illustration of distribution shifts for an MLLM (a) and performance degeneration and embedding shifts of the MLLM (b). An MLLM (LLaVA-v1.5-7B) receives arbitrary queries that might be visually and/or textually perturbed by unexpected noise. These distribution shifts result in performance drops, as shown in the middle bar plot. A visualization of intermediate layer representations of the MLLM on LLaVA-Bench-COCO and its variants indicates that MLLM fails to learn a proper level of invariance to generalize multimodal queries in the representation space.
Multimodal large language models (MLLMs).
Recent advances in MLLMs integrate a pre-trained language model with a vision encoder through visual instruction tuning liu2023visual ; dai2023instructblip . To be specific, let $X=(X_{v},X_{t})$ denote a multimodal input query consisting of visual and textual input, e.g., an image and a corresponding instruction or a question given that image, and $Y$ denote a desired response given the input query. An MLLM $f_{\theta}$ with parameter $\theta$ is trained to produce the desired response given an input query with a conditional autoregressive language modeling objective, i.e., $\arg\min_{\theta}\mathbb{E}_{X,Y}[\sum_{m=1}^{M}\log f_{\theta}(Y_{m}|X_{v},X_{t},Y_{<m})]$ for a sequence of $M$-length responses, where the visual input $X_{v}$ go through a visual encoder and projector modules to be converted as a sequence of tokens that have the same dimension as text embeddings and can be processed by an LLM backbone [^1] For simplicity, we will omit the visual encoder and projector in our learning objective at following sections.. After being trained, these models process a wide array of multimodal instructions to solve arbitrary visual question answering tasks lee2024vhelm .
Robustness problem in MLLMs.
Despite their impressive performance on standard benchmarks and their growing deployment in real-world applications li2024large ; li2024vision ; raza2025industrial , MLLMs remain vulnerable to input perturbations anonymous2024benchmarking ; cui2024robustness ; verma2024evaluating . For example, MLLMs undergo a systematic performance drop oh2025understanding when they encounter samples of superficial perturbations (e.g., varying brightness of image and typo in text) illustrated in Figure 1 (a). As shown in the bar plot of Figure 1 (b), LLaVA-v1.5-7B model undergoes severe performance degradation on LLaVA-Bench-COCO (LB-COCO; liu2023visual ) under the perturbations from visual input, textual input, and their joint (V, T, and J Pert.).
We posit that these vulnerabilities arise from the way MLLMs structure their internal representation space. In particular, inputs affected by perturbations are often embedded far from their intact (clean) counterparts, reflecting a distribution shift in the representation space that leads to poor generalization from an information-theoretic perspective oh2025understanding . The right side of Figure 1 (b) illustrates this phenomenon: using LLaVA-v1.5, we visualize representations of LB-COCO alongside its challenging variant, where the image and text inputs are perturbed. In this setting, semantically equivalent examples are mapped to distinct and distant regions in the latent space, suggesting a lack of invariance to superficial input variations, which is crucial for robustness to distribution shifts.
Motivated by this, our work aims to enhance the robustness of MLLMs by explicitly regularizing their internal representations, encouraging them to retain task-relevant information while discarding input-specific noise—thereby finding a good balance between invariance to low-level superficial features and sensitivity to high-level abstract features for better generalization.
Information bottleneck principle. The information bottleneck framework provides a principled approach to measure the quality of representations that are maximally predictive of a target variable while compressing redundant information from an input variable tishby2000information ; tishby2015deep . Numerous works have explored the use of IB training objective alemi2017deep , across computer vision luo2019significance ; Federici2020Learning , natural language processing mahabadi2021variational ; li2025calibrating , graph learning wu2020graph ; miao2022interpretable , and time-series modeling liu2024timex++ . These efforts are supported by theoretical insights suggesting that optimizing for the IB objective can reduce generalization error 8437679 ; kawaguchi2023does . However, most prior work focused on classification settings mahabadi2021variational ; li2025calibrating and/or relatively small-scale models alemi2017deep ; wang2021revisiting ; mahabadi2021variational . Although a recent study explored IB for MLLMs bai2025mitigating , the authors adopted IB training on a lightweight projector module while keeping the LLM backbone frozen. In contrast, our work is the first to investigate the IB framework for end-to-end training of large-scale autoregressive multimodal language models. Beyond shallow adaptations, we directly modify the internal structure of the LLM to promote IB-consistent behavior throughout the training process. We focus specifically on instruction tuning for MLLMs—which have become increasingly central to modern AI ecosystems but remain largely unexplored from the perspective of IB-based learning.
3 Method
3.1 Preliminary: Information Bottleneck As a Learning Objective
Let $X$ be a multimodal input query (e.g., image-text pair), $Y$ the desired output, and $Z=f(X)$ an intermediate representation extracted by the MLLM encoder $f(\cdot)$. The Information Bottleneck principle aims to learn representations that are maximally informative about the output $Y$ while being minimally informative about the input $X$. Formally, this is expressed as the optimization objective:
$\max_{f}\text{IB}_{f}(X,Y):=\underbrace{I(Z,Y)}_{\text{acquiring information from desired output}}-\beta\underbrace{I(Z,X)}_{\text{compressing input-specific redundant information}},$ (1)
where $I(\cdot,\cdot)$ denotes mutual information and $\beta$ is the trade-off coefficient. Minimizing $I(Z,X)$ encourages removing redundant or input-specific variations, while maximizing $I(Z,Y)$ ensures that the representation retains task-relevant signals necessary to predict the desired output.
In other words, the IB objective promotes representations that discard non-essential features tied to the input modality, while preserving those critical for solving the task. This property is particularly desirable for robust instruction tuning, where diverse multimodal inputs must be mapped to consistent, meaningful outputs under varied conditions (e.g., visual and textual perturbations). Despite its appeal, integrating the IB objective into MLLM training is highly non-trivial due to the intractability of mutual information estimation and the complexity of autoregressive and multimodal architectures.
3.2 Variational Inference for Information Bottleneck in MLLMs
Directly optimizing the IB objective is generally intractable, as it involves mutual information terms over unknown data distributions. In this work, we introduce a tractable variational bound on the IB objective, specifically tailored to the autoregressive and multimodal structure of MLLMs. We outline the key steps below and provide full derivations in the Appendix C.
We begin with the mutual information term $I(Z,X)$. Given the sequential nature of MLLMs, we decompose both the input $X=(X_{v},X_{t})$ and the latent representation $Z=(Z_{v},Z_{t})$ into visual and textual components. We can then derive the following upper bound for $I(Z,X)$:
$I(Z,X)$ $=\mathbb{E}_{x,z}[\log\frac{p(z|x)}{p(z)}]\leq\mathbb{E}_{x,z}[\log\frac{p(z|x)}{r(z)}]=\mathbb{E}_{x_{v},x_{t},z_{v},z_{t}}[\log\frac{p(z_{t}|x_{v},x_{t})p(z_{v}|x_{v})}{r(z_{v})r(z_{t})}]$
$=\mathbb{E}_{x_{v},x_{t}}[\mathbb{E}_{z_{t}|x_{v},x_{t}}[\mathbb{E}_{z_{v}|x_{v}}[\log\frac{p(z_{v}|x_{v})}{r(z_{v})}]]]+\mathbb{E}_{x_{v},x_{t}}[\mathbb{E}_{z_{v}|x_{v}}[\mathbb{E}_{z_{t}|x_{v},x_{t}}[\log\frac{p(z_{t}|x_{v},x_{t})}{r(z_{t})}]]]$
$=\mathbb{E}_{x_{v}}[D_{\rm KL}(p(z_{v}|x_{v})||r(z_{v}))]+\mathbb{E}_{x_{v},x_{t}}[D_{\rm KL}(p(z_{t}|x_{v},x_{t})||r(z_{t}))],$ (2)
where the first inequality holds given the non-negativity of Kullback-Leibler divergence (KLD), $D_{\rm KL}(r(z)||p(z))$, and $p(z_{v}|x_{v},x_{t})=p(z_{v}|x_{v})$ due to causal attention in MLLM. We introduce $r(z)=r(z_{v},z_{t})=r(z_{v})r(z_{t})$ as a factorizable variational approximation of the true prior $p(z)$.
Next, for the output-relevant term $I(Z,Y)$, we have the lower bound:
$I(Z,Y)=\mathbb{E}_{y,z}\left[\log\frac{p(y|z)}{p(y)}\right]\geq\mathbb{E}_{x,y,z}\left[\log q(y|z)\right]-\mathbb{E}_{y}[\log p(y)]\geq\mathbb{E}_{x,y}\left[\mathbb{E}_{z|x}\left[\log q(y|z)\right]\right],$ (3)
where we replace the true posterior $p(y|z)$ with a variational approximation $q(y|z)$ that will be parameterized by a model (will be elucidated in Section 3.3).
Finally, combining the lower bound of $I(Z,Y)$ and the upper bound of $I(Z,X)$ yields a variational lower bound for the IB objective as follows,
$\text{IB}(X,Y)\geq~{}~{}$ $\mathbb{E}_{x,y}\left[\mathbb{E}_{z|x}[\log q(y|z)]\right]$
$-\beta\;(\mathbb{E}_{x_{v}}\left[D_{\rm KL}(p(z_{v}|x_{v})||r(z_{v}))\right]+\mathbb{E}_{x_{v},x_{t}}\left[D_{\rm KL}(p(z_{t}|x_{v},x_{t})||r(z_{t}))\right]),$ (4)
In the next section, we elaborate on how we can implement this variational lower bound for MLLM instruction tuning in practice.
3.3 Vittle: A Practical Implementation of Visual Instruction Bottleneck Tuning
By using Monte Carlo approximation of expectations over data, Eq. (3.2) can be expressed as follows,
$\mathcal{L}_{\beta}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{z|x^{i}}[\log q(y^{i}|z)]-\beta\;(D_{\rm KL}(p(z_{v}|x^{i}_{v})||r(z_{v}))+D_{\rm KL}(p(z_{t}|x^{i}_{v},x^{i}_{t})||r(z_{t}))).$ (5)
To compute this empirical estimate of the IB lower bound, we need to model the posterior distributions, $p(z_{v}|x_{v})$ and $p(z_{t}|x_{v},x_{t})$, and prior distributions $r(z_{v})$ and $r(z_{t})$, of the MLLM’s inner representation $Z$. While in principle these distributions can take arbitrary forms, multivariate Gaussian distributions have been widely adopted in variational inference and probabilistic embedding literature graves2011practical ; KingmaWelling2014 ; Blei03042017 ; alemi2017deep ; oh2018modeling ; chun2021probabilistic due to their mathematical tractability and empirical effectiveness. By following this common standard, we set the posteriors and priors as Gaussian with diagonal covariance for $d$-dimensional variable, and will elucidate how exactly they are defined below.
Figure 2: Vittle architecture. We insert a learnable bottleneck layer $g_{\phi}=\{g_{\phi_{v}},g_{\phi_{t}}\}$ on top of $l$ blocks of LLM backbone (i.e., LLM-stem $f_{\theta^{l}}$) to estimate posterior distributions of token embeddings. After obtaining a sample per token $\{\tilde{z}_{v},\tilde{z}_{t}\}$ from posteriors, we interpolate it with a pre-bottlenecked token representation $\{z_{v},z_{t}\}$ and pass it through the remaining LLM blocks (i.e., LLM-head $f_{\theta^{l+}}$).
Posterior distributions.
As illustrated in Figure 2, we parameterize the posteriors $p(z_{v}|x_{v})$ and $p(z_{t}|x_{v},x_{t})$ using simple MLP blocks. Specifically, we introduce two non-linear projections, $g_{\phi_{v}},g_{\phi_{t}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{2d}$, which map each $d$-dimensional token embedding to the posterior Gaussian parameter vectors $\mu\in\mathbb{R}^{d}$ and $\sigma^{2}\in\mathbb{R}_{+}^{d}$ for the vision and language modalities, respectively. Given an intermediate $l$-th layer representation $(z_{v},z_{t})=f_{\theta^{l}}(x_{v},x_{t})$, we define:
$p(z_{v}|x_{v})=\mathcal{N}(z_{v};\mu_{v},\sigma_{v}^{2}\cdot I),\quad p(z_{t}|x_{v},x_{t})=\mathcal{N}(z_{t};\mu_{t},\sigma_{t}^{2}\cdot I),$
where $[\mu_{v},\sigma_{v}^{2}]=g_{\phi_{v}}(f_{\theta^{l}}(x_{v}))$ and $[\mu_{t},\sigma_{t}^{2}]=g_{\phi_{t}}(f_{\theta^{l}}(x_{v},x_{t}))$, with the mean and variance parameters split along output dimensions of MLP. These MLPs are applied position-wise in the same manner as Transformer’s feed-forward layers vaswani2017attention , producing token-wise variational posteriors. Now, we can sample from the posterior distributions of MLLM representation by $\tilde{z}_{v}\sim p(z_{v}|x_{v})$ and $\tilde{z}_{t}\sim p(z_{t}|x_{v},x_{t})$. Then, to strike a balance between invariance and sensitivity, we interpolate the original representation $z$ (pre-bottleneck) with its bottlenecked counterpart $\tilde{z}$ as $\hat{z}=(1-\alpha)z+\alpha\tilde{z}$. These representations are fed into the remaining layers to compute the predictive distribution over outputs, i.e., $q(y|z):=f_{\theta^{l+}}(y|\hat{z}_{v},\hat{z}_{t})$. While direct sampling introduces non-differentiability, we can enable the gradient flow using the reparameterization trick KingmaWelling2014 to sample $\tilde{z}$ via $\tilde{z}=\mu+\sigma\odot\epsilon$ with $\epsilon\sim\mathcal{N}(\mathbf{0},I)$ where $\mu$ and $\sigma$ are the outputs of the bottleneck MLP module given input $x$.
Prior distributions.
We consider two instantiations of the prior distribution for both $Z_{v}$ and $Z_{t}$: (1) a fixed standard Gaussian $\mathcal{N}(\mathbf{0},I)$, which is input-independent and enforces strong isotropy, and (2) a learnable Gaussian $\mathcal{N}(\mu_{\psi},\sigma^{2}_{\psi}\cdot I)$, where $\mu_{\psi}$ and $\sigma^{2}_{\psi}$ are two learnable vectors shared across samples. Each prior affects the formation of representations differently—the fixed prior imposes stronger regularization and robustness, while the learnable prior introduces additional flexibility by allowing the model to adapt to the instruction tuning distribution. We name the former Vittle (F) and the latter Vittle (L), and validate them altogether for all the evaluations in Section 4.
Overall objective and implementation.
The first term of $\mathcal{L}_{\beta}($Eq. (5)) can be easily computed through the standard cross-entropy, and our Gaussian instantiation of posterior and priors allows us to derive closed-form expressions of KLD terms that can be computed from simple arithmetic between $\mu$ and $\sigma^{2}$ parameters (See Appendix A.2). We set $\beta=\frac{0.1}{d}$ where $d$ is the hidden dimension of the MLLM, to normalize the KL regularization terms relative to the size of the latent dimension. The interpolation coefficient $\alpha$ in $\hat{z}=(1-\alpha)z+\alpha\tilde{z}$ increases progressively following a cosine schedule up to $0.5$. During inference, we consistently use an averaged representation $\hat{z}=(z+\tilde{z})/2$. The target layer to apply the bottleneck module can differ between visual and textual tokens, but we set $l=24$ for both modalities among 32 layers in a 7B-size LLM, i.e., top 25% layer, by default for simplicity (See Appendix B.1 for the ablation study). Figure 2 depicts the architecture overview.
3.4 Theoretical Justification
The learning objective of Vittle has an attractive theoretical interpretation that can support the improvement in robustness of Vittle. In this section, we first introduce a recently proposed information-theoretic measure of MLLM’s robustness under distribution shifts, effective mutual information difference (EMID oh2025understanding ), and show how Vittle can contribute to improving EMID.
Definition 3.1 (EMID).
Let $P_{\Theta}:\mathcal{X}\rightarrow{\mathcal{Y}}$ be an MLLM with parameters $\Theta$ that produces an output response $Y_{\Theta}$ given an input instruction $X$. For joint distributions $P_{XY}$ and $Q_{XY}$, effective mutual information difference of $P_{\Theta}$ over $P$ and $Q$ is defined as below,
$\text{EMID}(P_{XY},Q_{XY};{P_{\Theta}}):=[I(P_{XY_{\Theta}})-I(P_{XY})]-[I(Q_{XY_{\Theta}})-I(Q_{XY})].$ (6)
where $I(\cdot)$ denotes mutual information that measures the relevance between input instruction and response. A higher value of EMID indicates that MLLM $P_{\Theta}$ undergoes performance degeneration in the distribution $Q$ (test data) compared to $P$ (training data), so we want to achieve a lower value of it to ensure robustness. We now derive an upper bound for EMID (See Appendix D for the proof).
Proposition 3.2 (EMID upper bound).
Let $P_{\Theta}$ be an MLLM that maps $X=\{X_{v},X_{t}\}$ to $Z=\{Z_{v},Z_{t}\}$, and then sequentially maps $Z$ to $Y_{\Theta}$. Given joint distributions $P_{XY}=P_{X}\times P_{Y|X}$ and $Q_{XY}=Q_{X}\times Q_{Y|X}$ (resp. $P_{ZY}$ and $Q_{ZY}$), by assuming consistent conditionals over $Z_{v}|Z_{t}$, $Z_{t}|Z_{v}$, and $Y|X$ between $P$ and $Q$, we have an upper bound for $\text{EMID}(P_{XY},Q_{XY};{P_{\Theta}})$ as below,
$\hat{H}\big{(}D_{\rm JS}^{\frac{1}{2}}(P_{Z_{v}}||Q_{Z_{v}})+D_{\rm JS}^{\frac{1}{2}}(P_{Z_{t}}||Q_{Z_{t}})+\sqrt{\Delta_{X|Z}}\big{)}+|H(P_{Y_{\Theta}})-H(P_{Y})|+|H(Q_{Y_{\Theta}})-H(Q_{Y})|,$ (7)
where $H$ and $D^{\frac{1}{2}}_{\rm JS}$ indicate the entropy and square root of Jensen-Shannon divergence (JSD), respectively, $\Delta_{X|Z}:=\mathbb{E}_{z\sim P}[D_{\rm KL}(P_{X|z}||M_{X|z})]+\mathbb{E}_{z\sim Q}[D_{\rm KL}(Q_{X|z}||M_{X|z})]$ with a mixture distribution $M=\frac{P+Q}{2}$, and $\hat{H}:=\max_{x\in\mathcal{X}}[H(Q_{Y|x})+H(P_{Y_{\Theta}})]$. As we consider an optimization problem of $\Theta$, the terms, $H(P_{Y})$, $H(Q_{Y})$, and $\max H(Q_{Y|X})$, can be ignored from Eq. 7. We can also ignore $\sqrt{\Delta_{X|Z}}$ term because it cannot directly affect $Y_{\Theta}$ given the Markov assumption $X\rightarrow Z\rightarrow Y_{\Theta}$.
Implication. Vittle maximizes the variational lower bound of IB, which consists of (1) minimizing a standard negative log-likelihood term representing an expected risk, and (2) minimizing KLD terms to enforce posterior distributions close to prior distributions. By pursuing (1), MLLM $P_{\Theta}$ seeks a solution $\Theta$ that minimizes the expected risk and reduces its output entropy $H(P_{Y_{\Theta}})$ and $H(Q_{Y_{\Theta}})$ wen2024mitigating ; yang2024can ; groot2024overconfidence . Besides, it also reduces JSD between representation distributions $P_{Z}$ and $Q_{Z}$ by promoting all posterior samples to be laid near the pre-defined priors. In summary, reduced entropy and JSD terms induce lower EMID, which means that Vittle tries to achieve minimal difference between effective mutual information over training and evaluation distributions, while adapting to the in-distribution training set.
We show that Vittle indeed reduces JSD and EMID under distribution shifts in Table 4.2, and demonstrate in Section 4.2 that Vittle’s nice theoretical property is translated into consistent robustness gains under 30 distribution shift scenarios while maintaining in-distribution task performance.
4 Experiment
4.1 Setup
Model and implementation detail. We adopt LLaVA-v1.5 liu2024improved as our main baseline MLLM, where we set CLIP ViT-L/14-336px radford2021learning as a vision encoder, Vicuna-v1.5-7B vicuna2023 as an LLM, and a two-layer MLP as a projector. We follow the standard two-stage training of LLaVA liu2023visual , and replicate stage-1 for image-text alignment with the same configuration and dataset (LLaVA-pretrain-558k) of LLaVA-v1.5 liu2024improved . Then, on the LLaVA-mix-665k, we apply our Vittle objective. To validate the scalability and broad applicability, we also consider LLaVA-v1.5-13B and Prism-7B karamcheti2024prismatic . Refer to Appendix A for details and Appendix B for LLaVA-v1.5-13B and Prism-7B results, respectively.
Task, metric, and datasets. We evaluate instruction-tuned MLLMs with three representative tasks: (1) open-ended question answering, (2) object hallucination detection, and (3) closed-form question answering. All are formatted as a question answering (QA) with a single image input, where we use the average relative preference score measured by GPT-4o LLM judge zheng2023judging with three repeated runs for open-ended QA, while using exact matching accuracy for hallucination detection and closed-form QA. For open-ended QA tasks, we adopt four datasets: LB-COCO liu2023visual as a clean and typical dataset, and LLaVA-Bench in-the-wild (LB-Wild), LLaVA-Bench-Wilder (LB-Wilder), and WildVision-Bench (WV-Bench) as long-tail datasets. Then, we apply 27 types of image and text perturbations on LB-COCO samples [^2] See Appendix A.3 for a comprehensive summary of all perturbations and their generation processes. to yield 28 variants of perturbed LB-COCO (one of clean and nine of visual, textual, and joint perturbations, respectively). For object hallucination detection tasks, we adopt POPE li2023evaluating as a clean and typical dataset. Then, we generate nine variants of perturbed POPE with visual perturbations. Here, we consider the LB-COCO and POPE as in-distribution (ID) datasets because they are generated from MS-COCO samples that construct majorities of the instruction tuning set of modern MLLMs, including LLaVA. For closed-form QA, we adopt four representative datasets: ScienceQA lu2022learn , MMMU yue2024mmmu , MME liang2024survey , and MMStar chen2024are . In summary, we experiment with 45 datasets (31 of open-ended, 10 of object hallucination detection, and 4 of closed-form tasks).
4.2 Results
Figure 3: Object hallucination detection performance on POPE variants. We enumerate the hallucination detection accuracy of each method on nine versions of perturbed samples, and observe consistent gains by Vittle (highlighted by green numbers of relative improvement from baseline).
Figure 4: Open-ended QA performance on LB-COCO variants. We enumerate the relative preference score of responses from each model on 18 version of perturbed samples, and observe consistent gains by Vittle (especially for the Vittle (F)) on most of the textual (top), and joint (bottom) perturbations (results on visual perturbations are deferred to Appendix B).
Vittle improves robustness under input perturbations.
We first evaluate Vittle on object hallucination detection tasks with nine variants of POPE perturbed by visual corruptions in Figure 3. Although MLLMs trained with a standard objective and Vittle similarly suffer from perturbations, two instantiations of Vittle consistently outperform the standard objective. Interestingly, Vittle outperforms the baseline even in clean POPE (See Appendix B). We speculate that Vittle’s information control prevents the reliance on a partial feature of a single modality rahmanzadehgervi2024vision , e.g., textual feature, which is a common source of hallucination.
Figure 5: Evaluation under varying perturbation severity. Vittle achieves better performance, especially on severe perturbations.
Next, we present the validation on the open-ended QA task with 18 types of input perturbations, which are applied to visual and textual input independently or simultaneously in Figure 4. As we can see, Vittle greatly enhances performance in various perturbation datasets highlighted by green numbers that indicate the relative improvements of Vittle (F) compared to the baseline. Among the two variants of Vittle, Vittle (F) showcases better generalization under perturbations than Vittle (L), suggesting the benefits of a conservative zero-centered isotropic prior distribution to address a variety of subtle input perturbations. Next, we further explore Vittle’s robustness by evaluating varying perturbation severity. To be specific, we generate perturbations on three different degrees that determine how significantly the image or text would be changed. In Figure 5, we see that Vittle achieves better performance in general, where the margin becomes larger under severe perturbations. In summary, we observe a consistent gain by Vittle on the perturbed input setting across two tasks, which indicates that Vittle enhances the robustness to distribution shifts by pursuing the minimality of data representation.
Vittle improves generalization to long-tail distributions.
Not only subtle perturbations on input, but long-tail samples are also commonly encountered in many MLLM applications. In Table 4.2, we validate Vittle on three long-tail QA tasks constructed with real-world user queries. We see that Vittle also excels in generalizing long-tailed samples compared to the baseline. Interestingly, Vittle (L)–learnable prior–exhibits better performance compared with Vittle (F). We speculate that a learnable prior IB guides the model to learn a better sensitivity for high-level abstractions as well as an invariance to low-level noise by allowing additional flexibility to shape data-driven priors, yielding superior performance on tasks that require in-depth understanding of irregular queries.
Table 1: Performance comparison on long-tail open-ended QA tasks those contain queries that are quite different from typical training samples in terms of visual content and textual semantics.
Method LB-Wild LB-Wilder WV-Bench
Baseline 51.6 156.9 60.0
Vittle (L) 54.6 168.8 60.4
Vittle (F) 52.2 166.1 59.7
Table 2: Performance comparison on general benchmark datasets. These four multi-choice QA datasets require a higher level of multimodal understanding across multiple domains.
Method SciQA MMMU MME MMStar Avg.
Baseline 64.6 35.6 69.7 33.7 50.9
Vittle (L) 64.7 35.3 70.5 33.7 51.1
Vittle (F) 65.4 34.5 70.1 33.5 50.9
Vittle preserves competitive performance on general benchmarks.
Although the main focus of Vittle is to improve the model’s robustness under distribution shifts, securing the rich multimodal understanding capability and knowledge to diverse disciplines is also crucial as an essence of MLLM. To validate the multimodal understanding with an advanced level of knowledge, we evaluate each method on four representative closed-form QA benchmark datasets covering various fields. In Table 4.2, we observe that Vittle shows competitive performance with the standard approach, which implies that Vittle can also be used as a general-purpose learning objective.
Table 3: Comparison with weight-space compression methods. We compare Vittle with the LoRA and weight decay (WD) methods on LB-COCO and its perturbed variants.
Method Clean V Pert. T Pert. J Pert.
Baseline 77.8 73.4 72.2 62.3
LoRA 73.4 70.4 62.7 39.7
WD 74.1 72.1 73.0 59.5
Vittle (L) 76.7 73.9 73.0 62.7
Vittle (F) 76.1 74.2 74.1 64.4
Figure 6: Comparison with other objective functions. We report the average performance for all perturbations in POPE and LB-COCO.
Comparison with alternative learning approaches.
Note that the regularization forced by Vittle works on the representation space to penalize the amount of information encoded in the data representations. One of the natural alternatives is to regularize the model weight directly. In Table 4.2, we compare Vittle with LoRA hu2022lora and the weight decay method (WD) as instantiations of weight-space regularization, and the results suggest that explicit regularization on weight-space does not ensure a good balance between adaptability on in-distribution and robustness to distribution shifts. The other line of alternatives is information maximization during visual instruction tuning wang2025reconstructive ; zhou2025learning , which is the exact opposite of Vittle’s design principle. We compare two recent methods of this, ROSS wang2025reconstructive and LIT zhou2025learning , with Vittle on LB-COCO and POPE under perturbations. As shown in Figure 6, although these approaches are effective in improving object hallucination detection performances, they fail to achieve competitive performance on the open-ended QA task (See Appendix B for more results). This implies the non-trivial challenge of devising a general learning objective for MLLMs that can consistently improve robustness across diverse tasks, where we can see the promise of Vittle towards broadly applicable robust instruction tuning.
Qualitative analysis.
Figure 7 shows responses of clean queries and their visually or textually perturbed counterparts. Although the query before and after each perturbation conveys the same meaning and intention, LLaVA-v1.5 reveals volatility in its responses, whereas Vittle shows stable behavior by providing consistent responses, i.e., generating exactly the same response in the case of visual perturbation while keep focusing on the same object in a case of textual perturbation.
Figure 7: Case study on LB-COCO under perturbations. Although LLaVA-v1.5 produces a reasonable response for clean samples, the response and its quality vary under perturbations. Meanwhile, Vittle maintains the consistency for the responses.
Representation analysis.
We next see how Vittle shapes the representation space and how it affects robustness. In Table 4.2, we measure the average value of empirical JSD and EMID discussed in Section 3.4 over 27 perturbed variants of LB-COCO. Both JSD and EMID are computed between two distributions, clean and one of its perturbed versions, and then averaged over 27 clean-perturbed pairs (See Appendix A.4 for details). As our hypothesis, Vittle reduces distributional gaps, e.g., achieving smaller JSDs, between clean and perturbed samples in its representation space, thereby achieving a smaller EMID value that indicates better robustness. In Figure 8 (top), we further show PCA visualizations (in the same axis scale) for representations of LLaVA and Vittle on clean and image-text perturbed LB-COCO. We see that Vittle embeds the clean and semantically equivalent perturbed samples more closely. Moreover, the bottom panel shows that Vittle induces smaller cosine distances between clean and perturbed pairs in terms of the histogram and the average value in parentheses. These results indicate that our learning objective is indeed effective in structuring a better representation space that drives robustness.
Table 4: JSD and EMID evaluation on 27 LB-COCO variants.
Method JSD ($\downarrow$) EMID ($\downarrow$)
Baseline 0.068 0.026
Vittle (L) 0.048 0.021
Vittle (F) 0.047 0.025
Figure 8: PCA and pair-wise cosine distance of representations.
Table 5: Runtime comparison.
Method Tr.///it. (s) Tr. (h) Te. (s)
Baseline 7.363 11.06 0.1048
Vittle 9.482 13.36 0.1072
Cost analysis. Vittle introduces a lightweight bottleneck layer inside of LLM that slightly increases the total number of trainable parameters (by 1.5%). One may thus wonder how Vittle’s training and inference time is compared with a bottleneck-free baseline. In Table 5, we show the wall-clock training (per iter and total), and inference time per sample. Although Vittle increases the training time up to 20% compared with baseline, its inference time is almost identical to the original model, which is a reasonable amount of cost overhead given significant gains in terms of robustness.
5 Conclusion
This work provided the first investigation on the promise of information bottleneck from the context of MLLM instruction tuning to ensure robustness of MLLM under distribution shifts. We proposed a new theoretically-grounded visual instruction tuning method, Vittle, that injects a bottleneck layer inside the LLM to induce posterior samples of internal representations that encode useful information to produce valid responses while discarding other residual information from input queries. With negligible additional cost, Vittle is easily optimized with a variational lower bound of IB and shows consistent gains in robustness in 30 types of distribution shifts while also achieving strong performance on standard benchmarks, indicating that Vittle promotes a good balance between invariance and sensitivity during representation learning.
Acknowledgments and Disclosure of Funding
The authors thank Seongheon Park and Sean Xuefeng Du for their valuable suggestions and discussions that shaped the draft. Research is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation (NSF) Award No. IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, and Alfred P. Sloan Fellowship.
Appendix
Appendix A Extended Experiment Setup and Implementation Detail
A.1 Model and Training
In this work, we consider LLaVA-v1.5 liu2024improved as our target multimodal large language model (MLLM) with CLIP ViT-L/14-336px radford2021learning and Vicuna-v1.5-7B vicuna2023 as visual encoder and LLM backbone, respectively, and a two-layer MLP as projector (modality connector that maps features of the visual encoder into the text embedding space). Although all of the results presented in the main body of the paper were produced with LLaVA-v1.5-7B, we also experimented with LLaVA-v1.5-13B (with Vicua-v1.5-13B as the LLM backbone) to validate the scalability of our method, and consider Prism-7B karamcheti2024prismatic as an additional MLLM architecture to validate the broad applicability of Vittle. For fair comparison, all models are trained on the LLaVA-pretrain-558k and LLaVA-mix-665k datasets, consisting of a mixture of LAION schuhmann2021laion , CC sharma-etal-2018-conceptual , SBU ordonez2011im2text datasets with BLIP captions li2022blip and a mixture of LLaVA-instruct-158K and academic-task-oriented (V)QA datasets, respectively. Training configurations such as optimizer, learning rate, and batch size are summarized in Table A.1 and Table A.1. All training runs are conducted with eight A100-80GB GPUs with DeepSpeed ZeRO library. The shortest run takes roughly 11 hours, whereas the longest run takes about 14 hours. Now, we elaborate on the overall workflow of LLaVA and Prism below. Table 6: Hyperparameter list of Vittle training. We adopt exactly the same configurations with LLaVA-v1.5 liu2024improved for Stage 1 and 2. Config Stage1 Stage2 Global batch size 256 128 Batch size per GPU 32 16 Learning rate 1e-3 2e-5 Learning rate schedule Cosine decay w/ linear warmup Warmup ratio 0.03 Weight decay 0.0 Epoch 1 Optimizer AdamW Precision bf16 Table 7: Hyperparameter list of Prism-Vittle training. We adopt exactly the same configurations with Prism-DINOSigLIP-Controlled-7B karamcheti2024prismatic single stage training. Config Value Global batch size 128 Batch size per GPU 16 Learning rate 2e-5 Learning rate schedule cosine decay w/ linear-warmup Warmup ratio 0.03 Weight decay 0.1 Epoch 1 Optimizer AdamW Precision bf16
LLaVA is built with a pre-trained visual encoder that takes visual inputs, a pre-trained LLM backbone that takes text instructions, and a lightweight projector that maps features produced by the visual encoder into the text embedding space of LLM backbone so that the visually-grounded multimodal instruction input query can be processed by the LLM backbone. LLaVA undergoes a two-stage training: (1) The first stage takes into account modality alignment, where the projector is trained on image and corresponding instruction or caption with a conditional language modeling loss implemented by aggregating cross-entropy losses across response tokens, while the visual encoder and LLM backbone are frozen. (2) The second stage stands for the instruction tuning, where the projector and LLM backbone are jointly trained on multimodal instruction samples with the same conditional language modeling loss while the visual encoder is still frozen. This two-stage training has been considered a standard approach for developing MLLMs and is widely adopted chen2023shikra ; ye2023mplug ; tong2024cambrian .
Prism has a model architecture similar to LLaVA, but provides some valuable insight into the design of the MLLM training recipe, and we note two remarkable design choices of Prism that distinguish it from LLaVA: (1) incorporating multiple visual encoders rather than hosting a single visual encoder, and (2) reducing the two-stage alignment-then-instruction tuning into a single-stage instruction tuning. Note that different self-supervised visual representation learning induces features that have different strengths, and several works reveal the benefits of ensembling multiple different visual encoders to leverage complementary advantages tong2024eyes ; karamcheti2024prismatic ; shi2025eagle . Prism incorporates SigLIP zhai2023sigmoid and DINOv2 oquab2024dinov to enjoy both a robust global feature and a fine-grained local feature. Meanwhile, Karamcheti et al. karamcheti2024prismatic showed that the simplified single-stage training strategy can be a cost-effective alternative to the standard two-stage training.
To train these MLLMs, we consider five baseline approaches: (1) the standard full LLM fine-tuning with conditional language modeling loss, (2) parameter-efficient LoRA hu2022lora fine-tuning with the conditional language modeling loss, (3) conditional language modeling loss with weight decay regularization, (4) reconstructive visual instruction tuning (ROSS) wang2025reconstructive , and (5) learning to instruct (LIT) zhou2025learning . For the LoRA-based training configuration, we use the same one provided by the official LLaVA-v1.5 repository [^3] https://github.com/haotian-liu/LLaVA, and for the weight decay regularization, we select the regularization magnitude parameter among $\{0.1,0.01,0.001\}$ based on the POPE evaluation result. We now elucidate two competitive baseline methods, ROSS and LIT, in the following paragraphs. It is worth noting that these methods are designed to encode more (visual) information into the representation space, which is opposite to our Vittle’s design motivation that pursues a minimal sufficient representation for improving robustness to distribution shifts.
Reconstructive visual instruction tuning (ROSS) follows the two-stage training of LLaVA, but tries to reconstruct the visual inputs from the LLM backbone by adopting a regression or denoising learning objective in addition to language modeling loss during its second stage. By doing so, ROSS guides the MLLM to learn a much richer visual understanding, which is usually lacking in modern MLLMs tong2024eyes ; rahmanzadehgervi2024vision . The reconstruction target can be a raw RGB pixel value or the latent representation from an external visual encoder such as VQGAN esser2021taming or VAE KingmaWelling2014 , and ROSS requires an additional trainable module to reconstruct visual content, which is discarded during inference. We follow the training recipe from the official code repository [^4] https://github.com/Haochen-Wang409/ross to replicate ROSS-D-7B with the same visual encoder and LLM backbone to LLaVA-v1.5. For a fair comparison with LLaVA and Vittle, we train the ROSS with the same dataset (that of LLaVA-v1.5) for both training stages, while the original ROSS model was trained on a slightly larger dataset in the second stage.
Learning to Instruct (LIT) also focuses on the visual shortcomings of current MLLM and tries to improve the visual understanding capability of MLLM by incorporating an additional loss term that incentivizes the encoding of additional visual information. To be specific, while the cross-entropy loss in LLaVA’s conditional language modeling objective is aggregated through the response tokens only, LIT introduces an extra cross-entropy loss term, which is aggregated over the instruction (question) tokens only, thereby enforcing MLLM to learn to predict a proper textual instruction given an image. As LIT uses the same visual and language backbone model and training dataset as LLaVA-v1.5, we use the pre-trained checkpoint of LIT from Hugging Face [^5] https://huggingface.co/zhihanzhou/LIT-LLaVA-1.5-Vicuna-7B/tree/main for evaluation.
In Figure 6, we observe that while ROSS and LIT are somewhat effective in improving performance on object hallucination detection tasks with the aid of enhanced visual understanding capability, they significantly underperform Vittle and even the original LLaVA on the open-ended QA task under distribution shifts. This implies that pursuing more information encoding during visual instruction tuning may not result in better robustness to distribution shifts, but aiming to learn a minimal sufficient representation via Vittle can be a promising solution for this (See Table 11 for details).
A.2 Vittle Implementation Details
[…]
Then, we jointly train the LLM backbone, the projector, and this bottleneck layer together during the second stage of training with the objective function 5. As we assume a diagonal covariance Gaussian for the prior and posterior distributions, Kullback–Leibler divergence (KLD) between the prior $p$ and posterior $q$ can be easily expressed as below,
$D_{\rm KL}(q,p)=\frac{1}{2}\sum_{j=1}^{d}\big{(}\log\frac{\sigma^{2}_{p}[j]}{\sigma^{2}_{q}[j]}-1+\frac{(\mu_{p}[j]-\mu_{q}[j])^{2}}{\sigma^{2}_{p}[j]}+\frac{\sigma^{2}_{q}[j]}{\sigma_{p}^{2}[j]}\big{)}$ (8)
where $\mu_{\cdot}$ and $\sigma_{\cdot}$ denote $d$-dimensional distributional parameter vectors and $[j]$ indicates $j$-th element from the vectors. Vittle has two important hyperparameters: (1) target layer index $l$ for bottleneck application, and (2) posterior KLD regularization strength parameter $\beta$. After tuning across $l\in\{24,28,31\}$ and $\beta\in\{\frac{0.01}{d},\frac{0.05}{d},\frac{0.1}{d},\frac{0.2}{d},\frac{1.0}{d}\}$ where $d$ denotes the latent dimension of the LLM backbone, we set $l=24$ and $\beta=0.1/d$ based on the average performance of POPE and LB-COCO clean datasets. The interpolation coefficient $\alpha$ can also be tuned, but we found that increasing $\alpha$ beyond 0.5 hinders stable training and observing increased language modeling loss at the later parts of training progress. Figure 12 and Table 10 present the results of hyperparameter ablation study.
def reparam(mu, logvar):
std = (logvar / 2).exp()
batch_size, seq_len, hidden_dim = mu.shape
z = torch.randn(batch_size, seq_len, hidden_dim)
return mu + std * z
def forward(self, input_embeds, img_seq_len, a, **kwargs):
hidden_states = input_embeds
for l_idx, llm_layer in enumerate(self.llm.layers):
layer_outputs = llm_layer(hidden_states, **kwargs)
hidden_states = layer_outputs[0]
if l_idx == self.bottleneck_layer_idx:
# posterior inference
v_params = self.g_v(hidden_states[:,:i_seq_len,:])
t_params = self.g_t(hidden_states[:,i_seq_len:,:])
v_mean = v_params[:,:,:self.h_dim]
v_logvar = v_params[:,:,self.h_dim:]
t_mean = t_params[:,:,:self.h_dim]
t_logvar = t_params[:,:,self.h_dim:]
v_post = reparam(v_mean, v_logvar)
t_post = reparam(t_mean, t_logvar)
z_post = torch.cat((v_post, t_post))
# interpolation between original and bottlenecked
hidden_states = (1-a) * hidden_states + a * z_post
Listing 1: Forward pass of Vittle
def normalized_kld(mu, logvar, modality=None):
if modality is None:
# vittle (F) - fixed prior N(0,I)
kl_loss = -0.5 * (1+logvar-mu**2-logvar.exp()).mean()
else:
# vittle (L) - learnable prior
mu_pr, logvar_pr = self.l_prior[modality]
logvar_d = logvar-logvar_pr
scaled_mu_d = (mu-mu_pr).pow(2)/logvar_pr.exp()
var_ratio = logvar.exp()/logvar_pr.exp()
kl_loss = -0.5 * (1+logvar_d-scaled_mu_d-var_ratio).mean()
return kl_loss
def loss(self, logits, labels, v_mean, v_logvar, t_mean, t_logvar):
lm_loss = self.llm.loss_function(logits, labels)
if self.learnable_prior:
flag_v, flag_t = "v", "t"
else:
flag_v, flag_t = None, None
kld_v = self.normalized_kld(v_mean, v_logvar, flag_v)
kld_t = self.normalized_kld(t_mean, t_logvar, flag_t)
return lm_loss + self.beta * (kld_v + kld_t)
Listing 2: Training objective of Vittle
Figure 9: PyTorch-style pseudo code for the forward pass and training objective of Vittle
A.3 Downstream Task Benchmark Construction
Open-ended QA task.
One of the most representative applications of an MLLM is the generation of free-form responses given multimodal instruction queries. We consider LLaVA-Bench COCO (LB-COCO; liu2023visual ) as a typical in-distribution (ID) open-ended QA dataset, which is constructed from MS-COCO images lin2014microsoft with GPT-generated text queries that have 90 pairs of image and text. We then generate 27 variants of this LB-COCO by applying nine types of image perturbations, nine types of text perturbations, and nine types of image-text joint perturbations, to benchmark MLLMs’ robustness under various distribution shifts (which will be elaborated at the end of this section). Meanwhile, we also consider three datasets LLaVA-Bench in-the-wild (LB-Wild; liu2023visual ), LLaVA-Bench Wilder (LB-Wilder; li2024llavanextstrong ), and WildVision-Bench (WV-Bench; lu2024wildvision ) constructed by real-world web users’ image-text paired queries of 60, 128, and 500 samples, respectively, to validate models’ capability to address long-tailed queries in practice. This results in 31 different open-ended QA datasets in total: clean and 27 perturbed LB-COCO variants, and three long-tail datasets.
Object hallucination detection task.
[…]
Closed-form QA task.
There are numerous closed-form QA datasets that assess the internal knowledge of MLLMs from various perspectives. In this paper, we consider four representative datasets: ScienceQA lu2022learn , MMMU yue2024mmmu , MME liang2024survey , and MMStar chen2024are , which are designed to validate multimodal knowledge and understanding capability across various domains.
Distribution shift simulation.
[…]
Figure 10: Examples of visual perturbations.
Figure 11: Examples of textual perturbations.
Table 8: List of visual perturbations. We consider nine visual perturbations from four categories: (1) Blur, (2) Digital, (3) Weather, and (4) Noise, to validate the robustness of MLLMs under diverse types of visual perturbations.
Name Category
Defocus Blur Blur
Zoom Blur Blur
Contrast Digital
Brightness Weather
Fog Weather
Frost Weather
Gaussian Noise Noise
Shot Noise Noise
Speckle Noise Noise
Table 9: List of textual perturbations. We consider nine textual perturbations from three categories: (1) character-level, (2) word-level, and (3) sentence-level, to validate the robustness of MLLMs under diverse types of textual perturbations.
Name Category
Char Typo Character-level Perturbation
Char Delete Character-level Perturbation
Char Insert Character-level Perturbation
Word Swap Word-level Perturbation
Word Delete Word-level Perturbation
Word Insert Word-level Perturbation
Arabic Translation Sentence-level Perturbation
Greek Translation Sentence-level Perturbation
Hindi Translation Sentence-level Perturbation
A.4 Evaluation Details
Open-ended QA task.
Compared to multi-choice closed-form QA tasks that have a unique ground-truth answer per question, open-ended free-form generation-style QA tasks do not provide a single ground-truth answer. We follow the current standard evaluation paradigm, (M)LLM-as-a-Judge, that uses an external (usually more powerful) MLLM to gauge the quality of our target MLLM of interest via prompting. To be specific, for a given input query $x$, reference answer $y$, MLLM $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, and the judge model $r:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{Z}^{+}$, relative preference score is defined as, $\mathbb{E}_{x,y}[\frac{r(x,f_{\theta}(x))}{r(x,y)}]$.
For all of our open-ended QA evaluations, we used the same system prompt template provided by LLaVA authors [^7] https://github.com/haotian-liu/LLaVA/blob/main/llava/eval/table/rule.json, and we also adopted the MS-COCO annotation [^8] https://github.com/haotian-liu/LLaVA/blob/main/llava/eval/table/caps_boxes_coco2014_val_80.jsonl-based GPT-4 response [^9] https://github.com/haotian-liu/LLaVA/blob/main/playground/data/coco2014_val_qa_eval/qa90_gpt4_answer.jsonl and the gpt_answer [^10] https://huggingface.co/datasets/lmms-lab/LLaVA-Bench-Wilder released by LLaVA-NeXT authors as reference answers for LB-COCO variants and LB-Wilder, respectively. For LB-Wild and WV-Bench, we generated reference answers with GPT-4o.
Object hallucination detection and closed-form QA task.
[…]
Effective Mutual Information Difference (EMID) and Jensen-Shannon Divergence (JSD).
In addition to the evaluation with traditional metrics, we also consider the EMID and JSD-based evaluation, which was recently proposed as an information-theoretic approach to measure the robustness of MLLMs oh2025understanding . To compute the empirical estimates of MI, which is required for EMID computation, we use the CLUB estimator cheng2020club and reproduce the training and inference process of oh2025understanding by adopting image and text embeddings for the input image and text from CLIP-ViT-B/32 radford2021learning and XLM-RoBERTa-Base conneau2019unsupervised to replace $X_{v}$ and $X_{t}$, and also the text embeddings of XLM-RoBERTa-Base for responses $Y$ and $Y_{\Theta}$. To compute empirical estimates of JSD, we adopted the representation JSD estimator hoyos2023representation on top of the CLIP-ViT-B/32 and XLM-RoBERTa-Base, too.
Representation analysis.
Inspired by a recent work that reveals the importance of intermediate layer representation of the LLM backbone skean2025layer , we use the last input token embedding of the 24th layer (out of 32 layers in a 7B LLM backbone) for all experiments carried out in the representation space (Figure 1 (b) right, Figure 8, Figure 13, and the JSD computation in Table 4.2).
Appendix B Additional Results
B.1 Ablation Study
We first investigate two important hyperparameters for Vittle: (1) bottleneck layer index $l$ and (2) KLD regularization strength $\beta$, where we determined those parameter values based on the average performance on the clean POPE and LB-COCO, while not observing performance on perturbation datasets for fair model selection. We then further explore the impact of the interpolation coefficient $\alpha$, which plays a role in controlling the balance between the original representation and the bottleneck representation. Note that we could not conduct such an extensive search due to the computational burden of training 7B 13B scale models, so the hyperparameter values found here may not be optimal, and Vittle can achieve better results with further hyperparameter tuning.
Figure 12: Ablation study for the bottleneck layer index (left) and KLD regularization magnitude parameter $\beta$ (right).
[…]
For KLD regularization strength parameter ablation (Figure 12 right), we swept across $\{0.01,0.05,0.1,0.2,1.0\}$, and found that in the POPE dataset, strong regularization results in better performance, whereas it is not the case for LB-COCO. We choose $0.1/d$ as our default, which induces balanced clean-data performance on these two tasks.
Table 10: Ablation study for the representation interpolation coefficient $\alpha$ of the bottleneck layer. We observe that using the bottlenecked representation beyond the half portion of the total hinders the convergence of the language modeling loss.
Alpha POPE POPE V Shifts Avg. LB-COCO LB-COCO T Shifts Avg.
Baseline 86.98 84.12 77.8 72.3
0.1 87.22 84.20 77.9 73.1
0.25 87.34 84.47 75.6 73.1
0.5 87.71 84.90 76.7 73.0
0.75 Failed to converge
1 Failed to converge
We also explore the effect of the representation interpolation parameter $\alpha\in[0,1]$, which can be interpreted as a gating mechanism to control the information flow. As $\alpha$ approaches one, the later parts of the LLM backbone (LLM head in our notation) mainly use the information-penalized representation, while if $\alpha$ becomes smaller, the model strongly relies on the original representation. In Table 10, we observe that using too large values of $\alpha$ results in diverging language modeling loss, indicating that using a strongly penalized representation only cannot predict proper response tokens in sequence. Meanwhile, the larger value of $\alpha$ induces better POPE performance, whereas the trend is inconsistent in the LB-COCO data set, which is consistent with the observations from the previous ablation study in $\beta$.
B.2 Full Results of Pair-wise Cosine Distance
Figure 13: Pair-wise cosine distance of intermediate representations between clean LB-COCO and 27 versions of perturbed LB-COCO datasets. Vittle (F) consistently reduces the representation gap between the clean samples and their semantically equivalent perturbed ones.
We speculate that the performance degradation of MLLMs under perturbations originates from the representation discrepancy between clean and perturbed samples. That is, in the ideal case, a clean sample and its semantically equivalent perturbed sample should be closely mapped in the representation space, but current MLLMs did not shape the representation space in that way (see Figure 1 and Figure 8). In Figure 13, we provide the histograms of representation space pair-wise cosine distance between clean and perturbed examples in 27 types of perturbations. As we can see, Vittle (F) consistently mitigates the representation gap by reducing the pair-wise distance over diverse types of perturbations.
B.3 Full Results with LLaVA-v1.5-7B and LLaVA-v1.5-13B
[…]
Method POPE POPE V Pert. LB-COCO LB-COCO V Pert. LB-COCO T Pert. LB-COCO J Pert.
Baseline 86.98 84.12 77.8 73.4 72.2 62.3
LoRA 83.33 80.23 73.4 70.4 62.7 39.7
WD 87.22 83.97 74.1 72.1 73.0 59.5
ROSS 87.79 84.67 74.4 72.0 71.3 60.0
LIT 87.38 84.21 77.5 72.1 72.9 58.9
Vittle (L) 87.71 84.91 76.7 73.9 73.0 62.7
Vittle (F) 87.81 84.99 76.1 74.2 74.1 64.4
Table 12: Vittle on LLaVA-v1.5-13B model. We compare Vittle with the standard learning objective on LLaVA-v1.5-13B model that uses Vicuna-v1.5-13B as an LLM backbone. We set the bottleneck layer index $l=36$, interpolation coefficient $\alpha=0.5$, and bottleneck KLD regularization strength $\beta=\frac{0.1}{d}$. Vittle outperforms baseline on perturbed datasets while showing rivaling performance on the clean dataset.
Method POPE POPE V Pert. LB-COCO LB-COCO V Pert. LB-COCO T Pert. LB-COCO J Pert.
Baseline 87.14 84.02 76.9 73.5 73.8 64.6
Vittle (L) 87.22 84.85 76.6 74.5 74.0 65.4
Vittle (F) 87.32 84.65 76.8 74.2 73.9 65.3
Table 11 summarizes the overall results of our perturbation benchmarks on object hallucination detection (POPE) and open-ended QA tasks (LB-COCO). We note two findings here: (1) weight-space regularization methods, such as LoRA and WD failed to achieve reasonable performance; (2) although information maximization-based instruction tuning methods, such as ROSS and LIT, somewhat improve performance on POPE and its perturbation datasets, they greatly underperform Vittle, indicating a non-trivial challenge to design a versatile instruction tuning objective that can improve MLLMs on broad tasks. Meanwhile, we explore whether Vittle can be effective for a much larger model, e.g., a 13B-scale model. Table 12 shows that Vittle achieves consistent performance gains in object hallucination detection and open-ended QA tasks under distribution shifts, implying the scalability of our method.
Figure 14: Object hallucination detection performance on POPE perturbation datasets (top), and Open-ended QA performance on LB-COCO perturbation datasets (three below) of Prism-7B. We enumerate the accuracy for the object hallucination detection task and relative preference score for the open-ended QA task of each method on perturbed datasets, where we observe consistent performance gains by Vittle.
B.4 Applicability to Other MLLMs
We now investigate Vittle’s effectiveness on another recent MLLM, Prism-7B, beyond LLaVA. As noted in Section A.1, Prism has quite a different design principle than LLaVA with respect to the visual encoder and the training strategy, so it is suitable for investigating the versatility of Vittle between models. Table 13 shows summarized results on our perturbation benchmarks [^11] Due to resource constraints, we only explore Vittle (F) one of our prior distribution instantiations.. In object hallucination detection tasks, Vittle outperforms the standard cross-entropy only training baseline on clean and perturbed datasets. In open-ended QA tasks, Vittle consistently boosts performance in perturbation scenarios with large margins while maintaining performance on the clean dataset. The results of the perturbation-specific performance comparison are provided in Figure 14.
Table 13: Vittle on Prism-7B model. We compare Vittle (F) with the standard learning objective under the Prism-7B model training regime that adopts two visual encoders (DINOv2 and SigLIP) and the single-stage training rather than two-stage training with a single CLIP visual encoder. Vittle significantly improves perturbation-robustness compared with a naive learning objective.
Method POPE POPE V Pert. LB-COCO LB-COCO V Pert. LB-COCO T Pert. LB-COCO J Pert.
Baseline 87.54 85.29 79.4 75.3 71.9 53.8
Vittle (F) 88.11 85.52 79.0 76.2 75.4 63.2
Appendix C Derivation of Variational Bound for IB in MLLM
Here we provide a full derivation for the variational lower bound for IB. The derivation skeleton was mainly inspired by existing works 8253482 ; alemi2017deep . We begin with the mutual information term $I(Z,X)$. Given the sequential nature of MLLM, we decompose both the input $X=(X_{v},X_{t})$ and the latent representation $Z=(Z_{v},Z_{t})$ into visual and textual components. We can then derive the following upper bound for $I(Z,X)$:
$I(Z,X)$ $=\int p(x,z)\log\frac{p(x,z)}{p(x)p(z)}dxdz=\int p(x,z)\log\frac{p(z|x)}{p(z)}dxdz$
$=\int p(x,z)\log\frac{p(z|x)}{r(z)}dxdz-D_{\rm KL}(p(z)||r(z))$
$\leq\int p(x,z)\log\frac{p(z|x)}{r(z)}dxdz$
$=\int p(x_{v},x_{t},z_{v},z_{t})\log\frac{p(z_{t}|x_{v},x_{t})p(z_{v}|x_{v})}{r(z_{v})r(z_{t})}dx_{v}dx_{t}dz_{v}dz_{t}$
$=\int p(x_{v},x_{t})\int p(z_{t}|x_{v},x_{t})\int p(z_{v}|x_{v})\log\frac{p(z_{v}|x_{v})}{r(z_{v})}dx_{v}dx_{t}dz_{v}dz_{t}$
$+\int p(x_{v},x_{t})\int p(z_{v}|x_{v})\int p(z_{t}|x_{v},x_{t})\log\frac{p(z_{t}|x_{v},x_{t})}{r(z_{t})}dx_{v}dx_{t}dz_{v}dz_{t}$
$=\mathbb{E}_{x_{v}}[D_{\rm KL}(p(z_{v}|x_{v})||r(z_{v}))]+\mathbb{E}_{x_{v},x_{t}}[D_{\rm KL}(p(z_{t}|x_{v},x_{t})||r(z_{t}))],$ (9)
where the first inequality holds given the non-negativity of $D_{\rm KL}[r(z),p(z)]$ and $p(z_{v}|x_{v},x_{t})=p(z_{v}|x_{v})$ due to causal attention in MLLM. Here, we introduce $r(z)=r(z_{v},z_{t})=r(z_{v})r(z_{t})$ as a factorizable variational approximation of the true prior for the latent representation $p(z)$.
Next, for the output-relevant term $I(Z,Y)$, we have the lower bound:
$I(Z,Y)$ $=\int p(y,z)\log\frac{p(y,z)}{p(y)p(z)}dydz=\int p(y,z)\log\frac{p(y|z)}{p(y)}dydz$
$=\int p(y,z)\log q(y|z)dydz+D_{\rm KL}(p(y|z)||q(y|z))-\int p(y)\log p(y)dy$
[…]
$=\mathbb{E}_{x,y}\mathbb{E}_{z\mid x}\left[\log q(y|z)\right].$ (10)
where $p(x,y,z)=p(x)p(z|x)p(y|x)$ given the Markov chain assumption $Y\leftrightarrow X\leftrightarrow Z$, and $p(z|x,y)=p(z|x)$ holds given that the representation $Z$ can not directly depend on $Y$, and the entropy term of $y$, i.e., $\int-p(y)\log p(y)dy=H(Y)$, is ruled out due to its independence for optimization problem. Here, we replace the intractable $p(y|z)$ with a variational approximation $q(y|z)$ that will be parameterized by a model.
Finally, combining the lower bound of $I(Z,Y)$ and the upper bound of $I(Z,X)$ yields a variational lower bound for the IB objective as follows,
$\text{IB}(X,Y)\geq~{}~{}$ $\mathbb{E}_{x,y}\left[\mathbb{E}_{z|x}[\log q(y|z)]\right]$
$-\beta\;(\mathbb{E}_{x_{v}}\left[D_{\rm KL}(p(z_{v}|x_{v})||r(z_{v}))\right]+\mathbb{E}_{x_{v},x_{t}}\left[D_{\rm KL}(p(z_{t}|x_{v},x_{t})||r(z_{t}))\right]).$ (11)
Appendix D Missing Proof
D.1 Preliminary
We start by providing a definition of Mutual Information (MI) below.
Definition D.1 (Mutual Information (MI)).
For a joint distribution $P_{XY}$ over $\mathcal{X}\times\mathcal{Y}$, the mutual information with respect to $P_{XY}$ is defined as,
$I(P_{XY}):=\mathbb{E}_{x,y\sim P_{XY}}[\log\frac{P_{XY}(x,y)}{P_{X}(x)P_{Y}(y)}].$ (12)
If $X$ is an instruction and $Y$ is a corresponding response, we regard $I(P_{XY})$ as a relevance between the instruction and the response that can be seen as a possible quantification of instruction following capability of MLLMs. Effective MI is defined based on the MI as follows:
Definition D.2 (Effective Mutual Information (EMI) oh2025understanding ).
Given the joint distribution $P_{XY}$ and MLLM $P_{\Theta}$ parameterized with $\Theta$, the effective mutual information between the input and model response is defined as,
$\text{EMI}(P_{XY};P_{\Theta}):=I(P_{XY_{\Theta}})-I(P_{XY}),$ (13)
where $P_{XY_{\Theta}}$ denotes the joint distribution between the input $X$ and the output of the model $Y_{\Theta}$. Although the vanilla MI can also be used as a metric to evaluate models’ output response by $I(P_{XY_{\Theta}})$, the scale of it varies depending on the target data distribution which is undesired when our interest is to compare performance of model across multiple domains which can be addressed by EMI. Recall that we are ultimately interested in the performance difference of MLLMs across two different datasets, and this can be captured by the EMI difference (EMID) as follows:
Definition D.3 (EMID).
Let $P_{\Theta}:\mathcal{X}\rightarrow{\mathcal{Y}}$ be an MLLM with parameters $\Theta$ that produces an output response $Y_{\Theta}$ given an input instruction $X$. For joint distributions $P_{XY}$ and $Q_{XY}$, effective mutual information difference of $P_{\Theta}$ over $P$ and $Q$ is defined as below,
$\text{EMID}(P_{XY},Q_{XY};{P_{\Theta}}):=[I(P_{XY_{\Theta}})-I(P_{XY})]-[I(Q_{XY_{\Theta}})-I(Q_{XY})].$ (14)
By setting $P$ as an instruction tuning distribution (training data) and $Q$ as an arbitrary test time distribution (evaluation data), we prefer a model that has a smaller EMID value, which indicates better robustness under distribution shifts between $P$ and $Q$. Now, based on the original theorem provided by Oh et al. oh2025understanding , we are ready to derive a new upper bound for EMID tailored to our representation-centric visual instruction tuning setup.
D.2 A New Upper Bound for Effective Mutual Information Difference
We first review Lemma 1 of Shui et al. shui2022novel and its adapted version, a conditional entropy bound oh2025understanding as follows,
Lemma D.4 (Lemma 1 from Shui et al. shui2022novel ).
Let $Z\in\mathcal{Z}$ be the real-valued integrable random variable, and denoting two distributions on a common space $\mathcal{Z}$ by $P$ and $Q$ such that $Q$ is absolutely continuous w.r.t. $P$. If for any function $f$ and $\lambda\in\mathbb{R}$ such that $\mathbb{E}_{P}[\exp(\lambda(f(z)-\mathbb{E}_{P}(f(z))))]<\infty$, then we have:
$\begin{split}\lambda(\mathbb{E}_{z\sim Q}[f(z)]-\mathbb{E}_{z\sim P}[f(z)])&\leq D_{\rm KL}(Q||P)\\ &+\log\mathbb{E}_{z\sim P}[\exp(\lambda(f(z)-\mathbb{E}_{z\sim P}[f(z)]))]\end{split}$
Lemma D.5 (Conditional entropy bound oh2025understanding ).
Let $f(x):=H(Q_{Y|x})$ and $\hat{H}(Q_{Y|x}):=\max_{x\in\mathcal{X}}H(Q_{Y|x})$, given the marginal distributions $P_{X}$ and $Q_{X}$, and conditional distributions $P_{Y|X}$ and $Q_{Y|X}$, according to Lemma D.4, we have a conditional upper bound:
$\begin{split}\text{i})\ \ \mathbb{E}_{x\sim P}[H(Q_{Y|x})]-\mathbb{E}_{x\sim Q}[H(Q_{Y|x})]&\leq\hat{H}(Q_{Y|\mathbf{x}})\sqrt{2D_{\rm JS}(P_{X}||Q_{X})}.\end{split}$
Similarly, given the marginal distribution $P_{X}$ and $Q_{X}$, and an MLLM $P_{\Theta}$, let $f(x):=H(P_{\Theta}(\cdot|x))$ and $\hat{H}(P_{\Theta}):=\max_{x\in\mathcal{X}}H(P_{\Theta}(\cdot|x))$, then, according to Lemma D.4, we have another conditional upper bound:
$\begin{split}\text{ii})\ \ \mathbb{E}_{x\sim Q}[H(P_{\Theta}({\cdot|x})]-\mathbb{E}_{x\sim P}[H(P_{\Theta}(\cdot|x))]&\leq\hat{H}(P_{\Theta})\sqrt{2D_{\rm JS}(P_{X}||Q_{X})}.\end{split}$
Next, we should also need to formulate the relationship between JSD in the input space and JSD in the representation space, which is done through Lemma D.6.
Lemma D.6.
Let $f:\mathcal{X}\rightarrow\mathcal{Z}$ be an encoder that maps an input $X$ to a representation $Z$, for the input distributions $P_{X}$ and $Q_{X}$ and $f$-induced representation distribution $P_{Z}$ and $Q_{Z}$, we have an inequality below,
$\sqrt{2D_{\rm JS}(P_{X}||Q_{X})}\leq\sqrt{2D_{\rm JS}(P_{Z}||Q_{Z})}+\sqrt{\mathbb{E}_{z\sim P}[D_{\rm KL}(P_{X|z}||M_{X|z})]+\mathbb{E}_{z\sim Q}[D_{\rm KL}(Q_{X|z}||M_{X|z})]}$ (15)
where $M_{X|z}:=\frac{P_{X|z}\,+\,Q_{X|z}}{2}$.
Proof.
We start from the definition of JSD,
$D_{\rm JS}(P_{X}||Q_{X})=\frac{1}{2}D_{\rm KL}(P_{X}||M_{X})+\frac{1}{2}D_{\rm KL}(Q_{X}||M_{X}),\quad M_{X}=\frac{P_{X}+Q_{X}}{2}.$
By applying the chain rule of KLD under a deterministic map $X\rightarrow Z=f(X)$, we know that,
$D_{\rm KL}(P_{XZ}||M_{XZ})$ $=D_{\rm KL}(P_{Z}||M_{Z})+\int P_{Z}(z)D_{\rm KL}(P_{X|z}||M_{X|z})dz$
[…]
$\Leftrightarrow D_{\rm KL}(P_{X}||M_{X})$
Then, we have,
$D_{\rm JS}(P_{X}||Q_{X})=D_{\rm JS}(P_{Z}||Q_{Z})+\frac{1}{2}(\mathbb{E}_{z\sim P}[D_{\rm KL}(P_{X|z}||M_{X|z})]+\mathbb{E}_{z\sim Q}[D_{\rm KL}(Q_{X|z}||M_{X|z})]),$
which results in ineq. (15) by applying the triangular inequality after multiplying 2 on both sides. ∎
Now we derive a new upper bound for EMID, which is defined over the representation space rather than the previous one defined over the input space oh2025understanding in Proposition D.7.
Proposition D.7 (EMID upper bound).
Let $P_{\Theta}$ be an MLLM that maps $X=\{X_{v},X_{t}\}$ to $Z=\{Z_{v},Z_{t}\}$, and then sequentially maps $Z$ to $Y_{\Theta}$. Given joint distributions $P_{XY}=P_{X}\times P_{Y|X}$ and $Q_{XY}=Q_{X}\times Q_{Y|X}$, by assuming consistent conditional distributions over $Z_{v}|Z_{t}$, $Z_{t}|Z_{v}$, and $Y|X$ between $P$ and $Q$, we have an upper bound for $\text{EMID}(P_{XY},Q_{XY};{P_{\Theta}})$ as follow,
$\hat{H}\big{(}D_{\rm JS}^{\frac{1}{2}}(P_{Z_{v}}||Q_{Z_{v}})+D_{\rm JS}^{\frac{1}{2}}(P_{Z_{t}}||Q_{Z_{t}})+\sqrt{\Delta_{X|Z}}\big{)}+|H(P_{Y_{\Theta}})-H(P_{Y})|+|H(Q_{Y_{\Theta}})-H(Q_{Y})|,$ (16)
where $H$ and $D^{\frac{1}{2}}_{\rm JS}$ indicate the entropy and square root of Jensen-Shannon divergence (JSD), respectively, $\Delta_{X|Z}:=\mathbb{E}_{z\sim P}[D_{\rm KL}(P_{X|z}||M_{X|z})]+\mathbb{E}_{z\sim Q}[D_{\rm KL}(Q_{X|z}||M_{X|z})]$ with a mixture distribution $M=\frac{P+Q}{2}$, and $\hat{H}:=\max_{x\in\mathcal{X}}[H(Q_{Y|x})+H(P_{Y_{\Theta}})]$.
Proof.
Given the entropy-based definition of the mutual information, $I(P_{XY}):=H(P_{Y})-\mathbb{E}_{x\sim P}[H(P_{Y|x})]$, let $P_{Y_{\Theta}}=\mathbb{E}_{x\sim P}[P_{\Theta}(\cdot|x)]$ and $Q_{Y_{\Theta}}=\mathbb{E}_{x\sim Q}[P_{\Theta}(\cdot|x)]$, then, EMID can be expressed as follows,
$\text{EMID}(P_{XY},Q_{XY};P_{\Theta})$
$=\text{EMI}(P_{XY};P_{\Theta})-\text{EMI}(Q_{XY};P_{\Theta})$
$=(H(P_{Y_{\Theta}})-\mathbb{E}_{x\sim P}[H(P_{\Theta}(\cdot|x))]-H(P_{Y})+H(P_{Y|X}))$
$-(H(Q_{Y_{\Theta}})-\mathbb{E}_{x\sim Q}[H(P_{\Theta}(\cdot|x))]-H(Q_{Y})+H(Q_{Y|X}))$
$\leq(H(P_{Y|X})-H(Q_{Y|X}))+\big{(}\mathbb{E}_{x\sim Q}[H(P_{\Theta}(\cdot|x))]-\mathbb{E}_{x\sim P}[H(P_{\Theta}(\cdot|x))]\big{)}$
$+|H(P_{Y_{\Theta}})-H(P_{Y})+H(Q_{Y})-H(Q_{Y_{\Theta}})|$
$\leq\underline{(H(P_{Y|X})-H(Q_{Y|X}))}_{\text{ (A)}}+\underline{\big{(}\mathbb{E}_{x\sim Q}[H(P_{\Theta}(\cdot|x))]-\mathbb{E}_{x\sim P}[H(P_{\Theta}(\cdot|x))]\big{)}}_{\text{ (B)}}$
$+|H(P_{Y_{\Theta}})-H(P_{Y})|+|H(Q_{Y})-H(Q_{Y_{\Theta}})|.$ (17)
Moreover, we have the following inequality for $H(P_{Y|X})$ proposed by oh2025understanding ,
$H(P_{Y|X})-H(Q_{Y|X})\leq 4\mathbb{E}_{x\sim P}[D^{\frac{1}{4}}_{\rm JS}(P_{Y|x}||Q_{Y|x})]+\mathbb{E}_{x\sim P}[H(Q_{Y|x})]+\mathbb{E}_{x\sim Q}[H(Q_{Y|x})]$ (18)
By plugging inequalities in Lemma D.5 and ineq. (18) into the ineq. (D.2) to replace the terms (A) and (B), and given the consistent conditional distribution assumption for $Y|X$, i.e., $P_{Y|X}=Q_{Y|X}$, we have a much simpler upper bound as follows,
$\text{EMID}(P_{XY},Q_{XY};P_{\Theta})\leq\hat{H}\sqrt{2D_{\rm JS}(P_{X}||Q_{X})}+|H(P_{Y_{\Theta}})-H(P_{Y})|+|H(Q_{Y})-H(Q_{Y_{\Theta}})|,$
where $\hat{H}:=\max_{x\in\mathcal{X}}[H(Q_{Y|x})+H(P_{Y_{\Theta}})]$. Then, we can further replace the term $D_{\rm JS}(P_{X}||Q_{X})$ by using Lemma D.6 to get a bound defined by representation divergence as below,
$\text{EMID}(P_{XY},Q_{XY};P_{\Theta})$
$\leq\hat{H}(\sqrt{2D_{\rm JS}(P_{Z}||Q_{Z})}+\sqrt{\mathbb{E}_{z\sim P}[D_{\rm KL}(P_{X|z}||M_{X|z})]+\mathbb{E}_{z\sim Q}[D_{\rm KL}(Q_{X|z}||M_{X|z})]})$
$+|H(P_{Y_{\Theta}})-H(P_{Y})|+|H(Q_{Y})-H(Q_{Y_{\Theta}})|.$ (19)
Meanwhile, the chain rule of KLD and the definition of JSD with our consistent conditional distributions for $Z_{v}|Z_{t}$ and $Z_{t}|Z_{v}$, one can easily show that,
$2D_{\rm JS}(P_{Z_{v}Z_{t}}||Q_{Z_{v}Z_{t}})$ $=D_{\rm KL}(P_{Z_{v}Z_{t}}||M_{Z_{v}Z_{t}})+D_{\rm KL}(Q_{Z_{v}Z_{t}}||M_{Z_{v}Z_{t}})$
$=D_{\rm JS}(P_{Z_{v}}||Q_{Z_{v}})+D_{\rm JS}(P_{Z_{t}}||Q_{Z_{t}})$ (20)
Plugging Eq. 20 into ineq. (19) and applying the triangular inequality complete the proof. ∎
Appendix E Extended Literature Review
Compression for generalization.
There is a rich history in the machine learning field that connects compression of the model or its inner representation to generalization arora2018stronger , from the classical learning theory with Occam’s razor littlestone1986relating ; BLUMER1987377 and Minimal Description Length RISSANEN1978465 ; hinton1993keeping ; grunwald2005minimum to IB principle tishby2000information ; tishby2015deep , by suggesting models that provide minimal and simplest representation of data generalize better 8437679 ; hinton1993keeping ; kawaguchi2023does ; sefidgaran2023minimum analogy to human perception miller1956magical ; barlow1961possible ; zhaoping2025vision . Recently, Wilson wilson2025deep proposed a new generalization bound for contemporary large-scale models where the compressibility of a learning algorithm plays a key role in better generalization. According to that discussion, even the maximally flexible billion-scale model can have a small effective dimensionality (indicating the higher compressibility) by embracing soft inductive biases finzi2021residual , such as, a regularization term, to the learning problem. On top of these, IB-objective of Vittle can be understood as a soft inductive bias to seek a minimal sufficient representation that helps generalization for the challenging queries.
Robustness of fine-tuned foundation models.
Although large-scale pre-trained models have appealing generalization capability across diverse data instances from different domains, their fine-tuned counterparts usually hurts that strong generalization capability while being adapting on task-specific in-distribution samples kumar2022fine ; wortsman2022robust . This undesirable performance compromise between adaptation to in-distribution samples and generalization to samples from broad domains has spurred the community to work on robust fine-tuning of foundation models kumar2022fine ; wortsman2022robust ; lee2023surgical ; goyal2023finetune ; tian2023trainable ; oh2024towards ; hwang2024imagenet . This line of work addresses the adaptation-robustness trade-off by (1) introducing a regularization term tian2023trainable ; oh2024towards , (2) tweaking the training procedure kumar2022fine ; lee2023surgical , or (3) merging multiple models in the weight space wortsman2022robust ; oh2025dawin . However, almost all of the existing robust fine-tuning literature has focused on a discriminative model, such as CLIP radford2021learning , under classification setups. Although there are a few works on robust instruction tuning of MLLMs liu2024mitigating ; han2024towards , they do not specifically focus on improving robustness under diverse types of distribution shifts and propose a data-centric approach, i.e., expanding instruction tuning datasets in terms of quantity or diversity, that requires external MLLM-based data generation process and/or careful post-processing from humans. In this work, we take a representation-centric approach that modifies the learning objective of visual instruction tuning to efficiently enhance the robustness of MLLM under diverse distribution shifts (27 types in total).
Appendix F Limitation and Future Work
One of the potential concerns with IB is its reliance on the quality of $Y$, i.e., a gold response to given instruction, which is usually generated by another (M)LLM. As disclosed by Yeh et al. yeh2025challenges , the existing datasets for supervised fine-tuning are quite noisy, and we cannot ensure the advantage of IB on this noisy annotation setup. Moreover, IB alone does not guarantee the generalization of samples from completely different domains and may require additional domain information du2020learning ; li2022invariant ; ZHANG2023109737 . Investigating the potential of noisy annotation setups and domain generalization setups can be interesting future research problems. Meanwhile, a well-organized representation space by IB can be helpful for representation engineering or steering methods zou2023representation ; liu2025reducing that are also worth exploring for future work.
Appendix G Impact Statement
Multimodal large language models (MLLMs) today have many societal applications. This work tackles the robustness of MLLMs to distribution shifts between training and test time data. We observed a consistent improvement of our proposal Vittle in various types of visual and textual shifts, allowing users to trust the model more than before to safely use AI in a variety of environments. Moreover, although we focused on the robustness perspective in this work, improved invariance-sensitivity trade-off also benefits the fairness-discriminativeness trade-off, which is another crucial desideratum towards reliable AI. Meanwhile, even though its robustness to distribution shifts was improved, there are still potential misuse cases with MLLMs that can affect humanity by producing systematically biased outputs, given the existence of some adversarial data providers or attackers.